OpenStudy (anonymous):

How many combinations of 3 are there in 8 letter? (geometry)

OpenStudy (koikkara):

Using ABCDEFGH a) If no letter can be repeated: First letter of he 3 letter combo can be chosen in 8 ways. Second letter can be chosen only in 7 ways (because you can't repeat the 1st letter) Third letter can be chosen in 6 ways. So, total number of combinations = 8*7*6=336 b) If letters can be repeated, each letter of the combination can be chosen in 8 ways, so the number of combinations is 8*8*8=512

Directrix (directrix):

I think @rachelbetsyy is asking how many 3 letter combos there are in the 8-letter word Geometry. There are duplicates as someone has already stated.

OpenStudy (koikkara):

i think, now answer looks correct.

Directrix (directrix):

@Koikkara I don't know for sure but it looks as if you cranked out the number of permutations of 8 letters taken 3 at a time. Order matters in permutations but not in combinations.